OpenStudy (anonymous):

Find the point-slope form of (-3,1) ; m = -2) Btw, m is = to slope. So the slope or rate of change is -2.

myininaya (myininaya):

point slope form is y-y1=m(x-x1)

myininaya (myininaya):

just enter in your m and your (x1,y1)

OpenStudy (anonymous):

I didn't get that answer :/

myininaya (myininaya):

you can't replace the m with -2 ? and the (x1,y1) with (-3,1)?

OpenStudy (anonymous):

Wait I meant find the slope form.

myininaya (myininaya):

you don't mean point-slope form?

OpenStudy (anonymous):

Ohh nevermind. I get it. Thanks!

myininaya (myininaya):

there is point slope form which is y-y1=m(x-x1) there is slope-intercept form which is y=mx+b --- m is the slope (x1,y1) is any point given on the line b is the y-intercept